### v2

@@ 6.2.3 Visualization
CAT can enhance the self-attention of central nodes. In order to verify whether CAT enhances the central node’s self-attention and reduces the DE it suffers, we compare the final learned self-attention of all nodes before and after trimming. We take CAT-unsup as an example and visualize the self-attention improvement in Figure 13. It can be found that for the vast majority of nodes, the Trimmed graph obtained by CAT can make GAT pay more attention to the node itself and avoid the neighbors’ distraction; while very few nodes show a decrease in self-attention, it may be due to the fact that the node already has high self-attention before trimming and the neighbors received higher attention after trimming due to the reduction of competitors. Fortunately, this situation is very rare in the dataset and does not affect the overall discrimination ability of the model.
@@
Figure 13: Comparison of learned self-attention before and after graph trimming. The vertical coordinate represents the change of self-attention after graph trimming ad higher value represents the more significant enhancement of the central nodes’ self-attention.
@@
CAT can alleviate the discrimination ability degradation of GAT. In order to verify whether the model discrimination ability has been improved, we conduct dimensional reduction and visualization on the node representations learned by the model, and their corresponding Silhouette Coefficient (SC) is also calculated. As shown in Figure 14, we use TSNE to reduce the representation to two dimensions, where the higher SC represents the higher discrimination ability on different classes. We compared the original features of the data, the representations output by GAT, and the representations output by CAT variants, and compared their corresponding SC. The parameter settings with the highest node classification accuracy were chosen for both GAT and CAT as the represented result. We find that the representations obtained using GAT may involve a lower SC than the original features, where the discrimination ability degradation of GAT is manifested. In contrast, the representations obtained by CAT variants consistently get a higher SC than both the original features and the GAT representations, which implies that CAT can alleviate the discrimination ability degradation of GAT. It can be observed that the discrimination abilities of the three CAT variants are relatively close. Generally, CAT-sup demonstrates the highest discrimination capability, followed by CAT-semi, with CAT-unsup showing the lowest. Although the SC obtained by our method is still not high enough, it is enough to get some boost in rescuing the decline in discrimination ability caused by attention mechanism.
@@
Figure 14: The Silhouette Coefficient (SC) of learned node representation. The representations obtained by CAT consistently get a higher SC than both original features and GAT representations.
@@
CAT can map graph to a representation space that is closer to the ideal semantic space. When visualizing the dimensionality reduced representations, we found that although CAT could learn neither representation with a high SC nor easily separable boundary for heterophilic graphs, more nodes learned by CAT have a significant clustering tendency compared to GAT, which is manifested as more clustered structures in visualized figures. As shown in Figure 15, on the Chameleon dataset, CAT can identify more clusters than GAT such as the dark green and dark purple clusters. On the Cornell dataset, the node representations obtained by CAT bring nodes of the same class closer in the representation space such as the light green and dark purple clusters, which means the nodes locate closer to the cluster center and are easier to be distinguished from the clusters in other classes. As shown in Figure 16, for the same base GAT model, different CAT variants exhibit significantly distinct performances. As the base model, GATv3 learns the Semantic Cluster distribution with the lowest cluster cohesion and separation. There is an evident trend that with the more Class-level Semantic Cluster information, the variant of CAT is able to learn the clusters that are more compact and separable. It is apparently that the distances between clusters learned by CATv3-sup are maximized and nodes within a cluster is closest to the cluster center, while CATV3-unsup exhibits the opposite performance. This indicates the significance of Class-level Semantic Clusters.
@@
Figure 15: Visualization of embedding in the representation space learned by GAT and CAT-unsup.
Figure 16: Visualization of embedding of Wisconsin dataset learned by GATv3, GATv3-unsup, GATv3-semi and GATv3-sup.
@@ 7.1 Fundamental Hypothesis on Heterophilic Graphs
Improving the performance of GATs (GNNs as well) on the node classification task on heterophilic graphs is only a tangible task, the more fundamental issue behind improving the discrimination ability of the model is the assumption about the generation mechanism of the heterophilic graph. The Strong Homophily Hypothesis actually holds that connections between nodes are generated because they are sufficiently similar to each other, thus deriving an aggregation mechanism that relies on neighbors, which the heterophilic graphs obviously don’t hold. This can lead to an important question for heterophilic graphs, we depict this in Figure 17.
Question: what the fundamental hypothesis underlying heterophilic graphs be like? How to build better GNNs with heterophilic graphs, or how to build a brand-new graph representation learning model for heterophilic graphs, requires us to propose new inductive biases based on the heterophilic graphs’ generation mechanism [62]. This will be a challenging, landmark mission.
@@ 7.2 Limitations of CAT and Future Works
The lack of general hypothesis for heterophilic graphs. In this paper, we do not answer the question proposed in Section 7.1 directly in terms of the mechanism of data generation, but we anticipate that the model mechanism derived should be quite different from the existing Neighborhood Aggregation. Based on this insight, we tentatively offered a possible way, and have made a preliminary attempt on GAT: to make the model learn not to rely on all of its neighbors, but to concentrate on itself. Specifically, we employ causal inference as a tool to identify those neighbors that can help GAT concentrate more on the central node itself as much as possible. Our solution relies on the attention mechanism of GAT, which is not a generalized solution. How to derive a general heterophilic graph representation learning framework is an endeavor in the future.
@@
The lack of an effective way to learn optimal class-level Semantic Cluster. According to our Class-level Semantic Space Hypothesis, the ideal semantic space is supposed to be compact and separable where the Semantic Clusters can be located accurately. In Table 3, the comparison between CAT with different paradigms shows that supervised Semantic Clustering methods are more effective for learning class-level Semantic Clusters. However, considering the semi-supervised learning paradigm of node classification tasks, the Class-level Semantic Clustering Module adopting unsupervised or semi-supervised manner will be more reasonable. The challenge is the high dimensionality, sparsity, and low semantic expressiveness of node features in the dataset, posing a challenge for unsupervised and semi-supervised clustering methods to map graph nodes to an ideal semantic space. Therefore, it is challenging for unsupervised and semi-supervised clustering methods to map graph nodes to an ideal semantic space. More effective methods to learn better Class-level Semantic Space using less label information should be explored in the future, including unsupervised, semi-supervised and self-supervised learning methods. Training self-adaption modules is also an effective way to be considered.
@@
The lack of extension to Transformer-based graph learning methods. Since we only investigate the discrimination ability degradation of GNNs when meeting heterophilic graphs caused by the local attention-guided message passing mechanism, our framework is specifically designed for GATs. However, the Transformer [63], as a powerful neural network capable of handling sequential data, with its ability to capture global attention, can be transferred to graph learning tasks with a more powerful attention mechanism. Whether Graph Transformers [64, 65, 66] face the same challenges as GATs on heterophilic graphs, and how to extend the strategy behind this work to the Graph Transformer architecture are directions for future exploration.
Figure 17: Discussion on heterophilic graphs.
@@ 8 Conclusion
In order to cope with the significant degradation of node classification performance of GATs on heterophilic graphs, we propose a graph trimming method applied on GAT framework, Causally graph Attention network for Trimming heterophilic graph (CAT). Three representative GATs are employed as the base model and their discrimination ability can be significantly improved after trimming the graph by CAT. Specifically, we propose a new hypothesis for GATs on heterophilic graphs, Low Distraction and High Self-Attention, arguing that enabling the central node to concentrate on itself and reduce distraction from a portion of neighbors is an effective way to avoid degradation of discrimination ability of GATs. Based on this hypothesis, we use the attention distribution learned on the original graph as a signal to estimate the distraction caused by neighbors with the help of causal inference method, and then identify the Distraction Neighbors. Distraction Neighbors are further removed in the way of graph trimming, allowing the GAT model to achieve better node classification performance by maintaining self-attention. Compared with existing methods, our method eliminates the need to alter the model architecture of GAT or searching for more neighbors globally in the graph, but learns a new graph structure in the process of learning attention to get a better attention distribution. The experimental results show that our method is able to achieve a significant improvement in the performance of the node classification task for GATs on a total of seven heterophilic graphs in three different sizes. In addition, the framework of our method can be applied to any LAMP-driven models.

### v3

@@ 6.2.3 Visualization
CAT can enhance the self-attention of central nodes. To verify whether CAT enhances the central node’s self-attention and reduces the DE it suffers, we compare the final self-attention values learned by all nodes before and after trimming. We take CAT-unsup as an example and visualize the self-attention improvement after graph trimming in Figure 13. It can be observed that for the vast majority of nodes, the graph obtained by CAT can make the base GAT pay more attention to the nodes themselves and alleviate the neighbors’ distraction; while very few nodes exhibit decreased self-attention, possibly because the nodes already obtained high self-attention before trimming and their neighbors received more attention after trimming due to the reduction in the number of competitors. Fortunately, this situation is rare and does not affect the overall discrimination ability of the model.
@@
Figure 13: Comparison between the self-attention values learned before and after graph trimming. The vertical coordinate represents the change in self-attention after graph trimming and a higher value represents a more significant enhancement of the self-attention level.
@@
CAT can alleviate the degradation of discrimination ability exhibited by the GAT. To visualize whether the model’s discrimination ability is improved, we conduct dimensionality reduction on the learned node representations and calculate their corresponding silhouette coefficient (SC). As shown in Figure 14, we use t-SNE to reduce the representations to two dimensions, where a higher SC represents an enhanced ability to discriminate between different classes. We compare the original input features, the representations output by base GATs, the representations output by CAT variants, and their corresponding SCs. The parameter settings yielding the highest node classification accuracy are selected as the representative result. We observe that the representations obtained by GAT involve a lower SC compared to that of the original features, indicating the discrimination ability degradation exhibited by GAT. In contrast, the representations obtained by CAT variants consistently achieve the highest SC, which implies that CAT can alleviate the discrimination ability degradation. The discrimination abilities of the three CAT variants are relatively close. Generally, CAT-sup has the highest discrimination capability, followed by CAT-semi, with CAT-unsup performing the worst. Although the SC obtained by our method is not sufficiently high, it is adequate for achieving some improvement in mitigating the decrease in discrimination ability caused by LAMP.
@@
Figure 14: The silhouette coefficient (SC) of the learned node representations. The representations obtained by CAT consistently achieve higher SCs than do both the original features and GAT representations.
@@
CAT can embed graphs to a representation space approaching the ideal semantic space. We observe that more nodes learned by CAT have significant clustering tendencies compared to GAT, which is manifested as more clustered structures in visualized figures. As shown in Figure 15, on the Chameleon dataset, CAT can identify more clusters than GAT such as the dark green and dark purple clusters. On the Cornell dataset, the representations obtained by CAT bring nodes belonging to the same class closer in the representation space such as the light green and dark purple clusters, implying that the nodes are located closer to the cluster center and are easier to distinguish from the clusters in other classes. For different base models, Figure 16 exhibits a slight difference between GAT and GATv2, while GATv3 which is specifically designed for handling heterophilic graphs, learns more distinguishable representations. Nevertheless, CATv3-unsup is capable of learning more compact clusters compared to GATv3 such as the light green clusters. CATv3-semi and CATv3-sup can further learn superior representations. As shown in Figure 17, there is an evident trend that with more Class-level Semantic Cluster information, CAT variants can learn more compact and separable clusters. As the base model, GATv3 learns the Semantic Cluster distribution with the lowest cluster cohesion and separation. The distances between the clusters learned by CATv3-sup are maximized, and the nodes within a cluster are closest to the cluster center, while CATV3-unsup exhibits the opposite performance. This phenomenon highlights the significance of Class-level Semantic Clusters.
@@
Figure 15: Visualization of the embeddings in the representation space learned by GAT and CAT-unsup.
Figure 16: Visualization of the embeddings of the Texas dataset learned by different base GATs and variants of CATv3.
Figure 17: Visualization of the embeddings of the Wisconsin dataset learned by GATv3, GATv3-unsup, GATv3-semi and GATv3-sup.
@@ 7.1 Fundamental Hypothesis on Heterophilic Graphs
A fundamental issue behind improving the performance of GNNs on heterophilic graphs is the assumption about the generation mechanism of heterophilic graphs. The strong homophily hypothesis holds that connections between nodes are generated because they are sufficiently similar, thus deriving a neighboring aggregation mechanism, which the heterophilic graphs don’t hold. This raised an important question for heterophilic graphs, which we depict in Figure 18.
Question: What is the fundamental hypothesis underlying heterophilic graphs? How to build a brand-new graph representation learning mechanism for heterophilic graphs? It requires us to propose new inductive biases based on the generation mechanism of heterophilic graphs. This is a challenging, landmark mission.
@@ 7.2 Limitations of CAT and Future Works
The lack of general hypothesis for heterophilic graphs. In this paper, we hypothesized that the generation mechanism underlying heterophilic graphs will derive models different from the current neighboring aggregation models. Based on this insight, we offered a possible way, and have made a preliminary attempt on GATs: to make the node concentrate more on itself instead of relying excessively on all neighbors. Specifically, we employ causal inference methods to identify those neighbors that can help central nodes concentrate on themselves as much as possible. Our solution relies on the attention mechanism of GATs, which is not a generalized solution. Determining how to derive a general heterophilic graph representation learning framework is an endeavor for the future.
@@
The lack of an effective way to learn optimal class-level Semantic Cluster. According to our Class-level Semantic Space Hypothesis, the ideal semantic space is compact and separable. Considering the semi-supervised learning paradigm of node classification tasks, it is more reasonable for the Class-level Semantic Clustering Module to adopt an unsupervised or semi-supervised manner. The challenges concern high dimensionality, sparsity, and low semantic expressiveness of original node features. In the future, it is imperative to explore more effective methods for learning a better Class-level Semantic Space with less label information, including unsupervised, semi-supervised, and self-supervised learning methods. Training self-adaption modules is also explorable.
@@
The lack of extension for transformer-based graph learning methods. We only investigate the discrimination ability degradation of GNNs when meeting heterophilic graphs caused by the LAMP mechanism. However, the transformer [45], a neural network with a powerful global attention mechanism, can be transferred to graph learning tasks. Whether graph transformers [46] face the same challenges as GATs on heterophilic graphs, and how to extend the current strategy behind this work to the graph transformer architecture is worthy of future investigation.
To comprehensively and visually assess the proposed method, we applied SWOT analysis [47] in CAT. More future endeavors can be inferred from the SWOT matrix (Table 5), such as base model reinforcement and extension, and high-quality heterophilic graph benchmarks. The result clearly shows that the Distraction Effect and Distraction Neighbors identified in CAT have different practical implications in various scenarios and can be applied to analyzing real-world business datasets. For example, Distraction Neighbors may represent the different roles of friends in heterophilic social networks.
Table 5: SWOT matrix of CAT.
Internal Factors
Strengths Weakness
1. No need to alter the base GAT model. 2. No need to seek for more similar neighbors . 3. Plug-and-play and applicable to any LAMP-driven GATs. 4. Afford causal interpretation. 1. Performance relies on the discrimination ability of the base GAT model. 2. Performance relies on the label distribution of raw data.
External Factors
Opportunities Threats
1. High accuracy when label information is sufficient to uncover the category distribution. 2. High accuracy when an effective clustering method is implemented. 3. Explain the role of nodes in real-world business scenarios like social network user analysis. 1. Worse performance when label information is insufficient. 2. Worse performance when the adopted clustering method performs poorly. 3. Unavailable when the base GAT model fails to execute.
Figure 18: Discussion regarding heterophilic graphs.
@@ 8 Conclusion
To cope with the significant degradation of node classification performance exhibited by GATs on heterophilic graphs, we propose a Causal graph Attention network for Trimming heterophilic graphs (CAT). Three representative GATs are employed as the base model and their discrimination ability can be significantly improved after adopting CAT. Specifically, we propose a new hypothesis for GATs on heterophilic graphs, Low Distraction and High Self-Attention, which suggests enabling the central node to concentrate on itself and reduce distraction from neighbors. Based on this hypothesis, we leverage causal inference methods to estimate Distraction Effect and identify Distraction Neighbors. Distraction Neighbors are removed via graph trimming, allowing the base GAT model to achieve better node classification performance by maintaining self-attentions. Compared with existing methods, our method eliminates the need to alter the architecture of GATs or search for more neighbors globally; instead, it learns a new graph structure to obtain a better attention distribution. The experiments show that our method achieves significant performance improvements in node classification tasks on seven heterophilic graphs of three sizes. In addition, the framework of our method can be applied to any LAMP-driven model.